OpenStudy (anonymous):

How did (e^x-e^-x)/(e^x+e^-x) become (e^2x-1)/ (e^2x+1) ??

OpenStudy (asnaseer):

that is done by multiplying the numerator and denominator by \(e^x\)

OpenStudy (asnaseer):

\[\frac{e^x-e^{-x}}{e^x+e^{-x}}=\frac{e^x-e^{-x}}{e^x+e^{-x}}*\frac{e^x}{e^x}\]\[=\frac{e^x(e^x-e^{-x})}{e^x(e^x+e^{-x})}=\frac{e^{2x}-1}{e^{2x}+1}\]

OpenStudy (anonymous):

Now I get it!! Thanks! Last question, why do I need to multiply it to e^x

OpenStudy (asnaseer):

I assume this is related to your last question on \(\tanh{x}=\frac{1}{2}\)?

OpenStudy (anonymous):

Yes

OpenStudy (asnaseer):

ok, so you know that:\[\tanh{x}=\frac{e^x-e^{-x}}{e^x+e^{-x}}\] correct?

OpenStudy (anonymous):

Yes

OpenStudy (asnaseer):

we are therefore trying to solve:\[\tanh{x}=\frac{e^x-e^{-x}}{e^x+e^{-x}}=\frac{1}{2}\]

OpenStudy (asnaseer):

one way of doing that is to first simplify the left-hand-side

OpenStudy (asnaseer):

so that you are left with a fraction involving just \(e^{2x}\)

OpenStudy (asnaseer):

you saw how that was done, so we then end up with:\[\frac{e^{2x}-1}{e^{2x}+1}=\frac{1}{2}\]now you can just cross multiply and solve for \(e^{2x}\)

OpenStudy (asnaseer):

does that make sense?

OpenStudy (anonymous):

Yes! Thanks a lot!!

OpenStudy (asnaseer):

yw :)